# Fold catastrophe in breaking waves

### v1

@@ 3.2 Intrinsic timescale
We rely on the following simple physical picture for the dynamics of waves near breaking, referring to non-breaking steep-to-maximally-steep waves and breaking waves. Imagine tracking a special moving point $Q$ on the wave surface, such as the energetic hotspot (BoettgerJFM2023), and the associated speed $C=C_{e}$. Moving with that point, the fluid beneath the surface appears to flow more slowly than observed from a fixed frame, with a small relative velocity $U-C_{e}$. This motivates the definition of the small parameter
@@
The velocity gradient $U_{z}=\partial_{z}u|_{Q}$ measured at the moving point $Q$ governs the rate of change of the dynamics near breaking. As a focusing wave steepens, $U_{z}$ increases in magnitude, indicating that vertical motion transports significant horizontal momentum toward the surface. This process amplifies the horizontal fluid motion and enhances the convergence of kinetic energy near the crest during wave steepening. Such a concentration of horizontal momentum and energy flux ultimately leads the wave to break (BoettgerJFM2023; BoettgerPRF2024). This trend is observed in numerical simulations of unsteady breaking and non-breaking wave groups (BarthelemyJFM2018) (see bottom-left panel of Fig. 1). In the simulations, the velocity gradient $U_{z}$ increases as the waves steepen, reaching a maximum at breaking, where the numerical solver stopped because it was not designed to follow the post-breaking phase. The relavance of velocity gradients to wave breaking was first emphasized by PhillipsBanner1974, who showed that thin shear layers induced by wind can trigger breaking at reduced amplitudes (BannerPhillips1974).
Following Rovelli1991TimeHypothesis; rovelli2004quantum; rovelli2018order, we define the dimensionless relational fast time $\tau$ [^2] Time, in Rovelli’s relational framework, measures change, not an absolute flow. Temporal order arises only through correlations between observables. The surface elevation of a wave changes with respect to the reading of some other observable (clock), which in our case we choose as related to the vertical velocity gradient. that measures the relevant changes of the dynamics in relation to steepening as observed at a moving point $Q$ on the surface:
@@
where $t_{0}$ is a convenient initial time after which $U_{z}\geq 0$, ensuring that $\tau$ increases monotonically (see bottom-right panel of Fig. 1, note the smoothness of $\tau(t)$ as a result of the integration). This definition applies to waves propagating from left to right along the $x$-axis, where $U_{z}>0$ near the crest of steep waves. [^3] For waves propagating in the opposite direction (right to left), $U_{z}<0$ and one sets $\frac{d\tau}{dt}=-U_{z}$ to maintain a monotonically increasing $\tau$. As waves steepen, the horizontal velocity gradient $U_{z}$ measured at $Q$ grows, eventually becoming very large and formally diverging at the onset of breaking, for instance during the formation of jets associated with spilling. This motivates defining near breaking, the small parameter
@@
Both parameters $\beta,\epsilon$ are small near breaking, and $\beta\sim O(\epsilon)$, estimated from numerical simulations (BarthelemyJFM2018). For very steep wave crests near the breaking inception, where $U_{z}\gg 1$ ($\epsilon\ll 1$), our analysis indicates that the focusing parameter $A$ in Eq. (17) is bounded and of order $O(1)$. Near breaking, this implies that $\eta_{xx}$, proportional to surface curvature, grows as fast as $O(\epsilon^{-2})$ to keep $A$ finite. Here, the notation $O(\epsilon^{n})$ denotes terms of order $\epsilon^{n}$ or higher.
@@
We demonstrate that this timescale separation occurs when the special moving point $Q$ is chosen to coincide with the energetic hotspot (BoettgerJFM2023; BoettgerPRF2024) (for details, see Appendix 14). It captures the essential dynamics near the onset of breaking and it reveals a robust fold geometry that guides the evolution of steep waves near breaking.
Our analysis is grounded in first principles and supported by numerical simulations of nonlinear wave groups by BarthelemyJFM2018. The same separation could not be conclusively established when following other surface points, such as the inflection point (McAllister2023).
Figure 1: Top panel: snapshots of the (black) surface elevation $\eta(x,t)$ and (red) kinetic energy $\tilde{k}_{e}(x,t)$ of the largest non-breaking wave (labeled 22) in the simulated right-going wave group $C3N5A0.514$ (BarthelemyJFM2018) at the pre-focusing ($P$), focusing ($F$), and post-focusing ($R$) instants, in the middle top panel of Fig. 3. The energetic hotspot (red square) is located at the front of the wave before focusing and moves to the back after passing through the crest (blue square). The inflection point is also depicted (blue diamond). Bottom panel: (left) time profiles of the horizontal velocity gradient $U_{z}$ tracked at the energetic hotspot of the four major successive waves of the simulated right-going group $C3N5A0.514$ (BarthelemyJFM2018). These are labeled as $0,1,2,3$ with varying thickness to denote the increasing amplitude, and wave $3$ is breaking; (right) relational fast time $\tau$ as function of the absolute time $t$.
@@ 4 Following the energetic hotspot: the slow-fast model
evaluated at the free surface $z=\eta(x,t)$, where $k_{e}=(u^{2}+w^{2})/2$ is the kinetic energy density of the fluid (see Appendix 12.4). Drawing on BoettgerJFM2023, we track the most energetic point $Q$ on the wave surface, defined as the location where $\tilde{k}_{e}$ attains its maximum, that is where the spatial kinetic energy gradient $\partial_{x}\tilde{k}_{e}=0$. This energetic point can be regarded as the crest of the kinetic energy surface  $\tilde{k}_{e}(x,t)$. The speed $C_{e}$ of the energetic hotspot reduces to the group velocity for linear wave packets.
To highlight the relating motion of the wave surface crest and energetic point, or hotspot, we consider the simulated breaking wave group $C3N5A0.514$ reported in BarthelemyJFM2018. The four major successive waves of the group are labeled in order of increasing amplitude, with wave $3$ undergoing breaking. The evolution of the second-largest wave (wave $2$) and its kinetic energy surface is shown in the top panel of Fig. 1. As the wave grows in amplitude (point $P$, pre-focusing phase), the energy crest is located on the front face of the wave, slightly behind the surface crest. At focus (point $F$), where the wave attains its maximum height, the surface crest and the energetic point coincide at the same location. After focus (point $R$), the energetic point shifts to the back face of the wave, behind the surface crest. The other waves in the group have similar evolution, except wave $3$ that breaks near the focusing point (see (BarthelemyJFM2018) for more details).
During the growth phase, the surface crest slows down (FedeleJFMcrestspeed2020; Banner2014CrestSlowdown; FedeleEPL2014), while the energetic point speeds up. However, the acceleration of the energetic point is insufficient to prevent the surface crest from overtaking it. After focus, the surface crest accelerates, whereas the energetic hotspot decelerates. The corresponding crest and energetic point speeds are shown in the bottom-right panel of Fig. 1.
@@ 4.1 The slow-fast model
The energy gradient $\partial_{x}\tilde{k}_{e}$ vanishes at the energetic hotspot $Q$. Thus, Eq. (69) in Appendix 12.4) yields
@@
In Appendix 14, we proved that a slow-fast regime occurs near breaking around the maximum $A=1+O(\epsilon)$, and along the path $\Gamma_{2}$ in the $(m,s)$ plane where $m_{2}\sim(m+s)=0$ (see Eq. (35)). This is where the vertical transport of horizontal momentum density dominates the horizontal advection during focusing when a wave reaches its maximum height, or breaks (see Eq. (16)). Numerical simulations confirm this trend, as illustrated in Figure 4, and is further discussed in section 4.4.
@@
Then, $dC_{e}/d\tau=dC_{r}/d\tau\sim O(\epsilon^{3/2})$. Using these scalings, in Appendix 14.1 a closed slow-fast system is derived in the variables $(m,A)$ (see Eq. (108)),
@@
Such slow-fast timescale separation is observed in numerical simulations of nonlinear wave groups (BarthelemyJFM2018; Boettger2024). The top panels of Fig. 2 depict the time profiles of the wave parameters $m$, $s$ and $R=m+s$ tracked at the energetic hotspot of the four major successive waves of the simulated right-going group $C3N5A0.514$ (BarthelemyJFM2018). These are labeled in order of increasing amplitude, with wave $3$ breaking. Near breaking, $R=m+s$ varies adiabatically, more slowly than the surface slope $m$ and the kinematic parameter $s$ over the fast time $\tau$. Similarly, the focusing parameter $A$ and $\beta=C_{e}/U-1$ are also slow variables (see bottom panels of Fig. 2. Thus, the slow variable $A$ acts as an adiabatic forcing for the fast geometric variable $m$.
@@
Figure 2: Top panels: time profiles of the wave parameters $m$, $s$ and $R=m+s$ tracked at the energetic hotspot of the four major successive waves of the simulated right-going group $C3N5A0.514$ (BarthelemyJFM2018). Bottom panels: same for the slow parameters AA and $\beta=C_{e}/U-1$ tracked at the energetic hotspot. Waves are labeled as $0,1,2,3$ with varying thickness to denote the increasing amplitude, and with wave $3$ breaking. Characteristic slopes are also shown (red lines) and quantified in terms of $\epsilon=1/U_{z}\approx 0.1$.
Figure 3: Fold catastrophe for inviscid and irrotational waves in deep water. (Left) fold boundary $\partial\mathcal{F}_{0}$ in the $(m,A)$ plane and trajectories (solid lines) of the four major successive waves of the simulated right-going wave group $C3N5A0.514$ (BarthelemyJFM2018). These are labeled as $0,1,2,3$ at the beginning (early time) of the trajectories, and depicted with increasing thickness. The thickest (dark red) line indicates the largest breaking wave. Points $P$, $F$, and $R$ mark the pre-focusing, focusing, and post-focusing phases of the largest non-breaking wave ($2$) (see Fig.1). Theoretical predictions (dashed lines) from the slow-fast model are also depicted. The path of a linear packet is also reported (amplitude $\mu=ka=0.1$ and spectral bandwidth $\nu=0.3$). (Left) the desymmetrized half-domain $m\leq 0$.
@@ 4.2 Fold Catastrophe and Breaking Onset
The fold boundary $\partial\mathcal{F}_{0}$ is shown in the left panel of Fig. 3 and the stable and unstable manifolds of equilibria are depicted with arrows. Equilibria with negative (positive) slope are stable (unstable). These equilibria coalesce at the tip $O$ at $(A=1,m=0)$, forming a marginal state with slope $m=0$. Thus, $\partial\mathcal{F}_{0}$ is normally hyperbolic, except at $O$. This boundary is also symmetric with respect to $m$, reflecting the parity symmetry of the Euler equations in deep water, which allows the same wave to propagate in either direction. In such cases, only the horizontal velocity $U$, its gradient $U_{z}$ and the hotspot speed $C_{e}$ change sign, leading to a corresponding sign change in both $s$ and the slope $m$, except the focusing parameter $A$. The symmetry can be removed by applying symmetry reduction techniques described in ChaosBook (see section 4.9 and right panel of Fig. 3 that displays the desymmetrized domain).
The point $O$ is the marginal point of a fold catastrophe (Thom1975; zeeman1976catastrophe; arnold1984catastrophe). Indeed, we can rewrite the surface slope equation as the gradient flow
@@
defines a cubic fold surface (Thom1975; arnold1984catastrophe).
The right panel of Fig. 3 illustrates the trajectories (blue lines) in the plane $(m,A)$ of the four major successive waves in the simulated right-going wave group $C3N5A0.514$ reported by BarthelemyJFM2018. These are labeled as $0,1,2,3$ at the beginning (early time) of the trajectories. Line thickness increases with wave amplitude, with the thickest trajectory corresponding to the breaking wave ($3$). Points $P$, $Q$, and $R$ denote the pre-focusing, focusing, and post-focusing phases of the largest non-breaking wave ($2$), whose profiles are shown in Fig. 1. For comparison, the trajectory of a linear Gaussian packet is also included (wave steepness $\mu=ka=0.1$, spectral bandwidth $\nu=0.3$).
All trajectories of the four nonlinear waves, during their growing phase, are attracted toward the point $O$ along the stable manifold of the negative-slope equilibria, and the energetic hotspot lies just behind the wave crest. They follow the parabolic arc toward the point $O$, where the wave crest and the energetic hotspot coincide. The breaking wave ($3$) is the only one that crosses the boundary near $O$, thereby entering its interior and undergoing finite-time blowup, indicating the onset of breaking. The numerical simulations, however, could not resolve the dynamics in the immediate vicinity of, or beyond, breaking. In contrast, the trajectories of the other non-breaking waves are repelled along the unstable manifold of the positive-slope equilibria, leading to the defocusing phase of the crests. A similar behavior is observed in other simulated breaking wave groups (BarthelemyJFM2018).

### v2

@@ 3.2 Intrinsic timescale
We rely on the following simple physical picture for the dynamics of waves near breaking, referring to non-breaking steep-to-maximally-steep waves and breaking waves. Imagine tracking a special moving point $Q$ on the wave surface, such as the energetic hotspot (Boettger et al., 2023), and the associated speed $C=C_{e}$. Moving with that point, the fluid beneath the surface appears to flow more slowly than observed from a fixed frame, with a small relative velocity $U-C_{e}$. This motivates the definition of the small parameter
@@
The velocity gradient $U_{z}=\partial_{z}u|_{Q}$ measured at the moving point $Q$ governs the rate of change of the dynamics near breaking. As a focusing wave steepens, $U_{z}$ increases in magnitude, indicating that vertical motion transports significant horizontal momentum toward the surface. This process amplifies the horizontal fluid motion and enhances the convergence of kinetic energy near the crest during wave steepening. Such a concentration of horizontal momentum and energy flux ultimately leads the wave to break (Boettger et al., 2023, 2024). This trend is observed in numerical simulations of unsteady breaking and non-breaking wave groups (Barthelemy et al., 2018) (see bottom-left panel of Fig. 1). In the simulations, the velocity gradient $U_{z}$ increases as the waves steepen, reaching a maximum at breaking, where the numerical solver stopped because it was not designed to follow the post-breaking phase. The relavance of velocity gradients to wave breaking was first emphasized by Phillips and Banner (1974), who showed that thin shear layers induced by wind can trigger breaking at reduced amplitudes (Banner and Phillips, 1974).
Following Rovelli (1991, 2004, 2018), we define the dimensionless relational fast time $\tau$ [^2] Time, in Rovelli’s relational framework, measures change, not an absolute flow. Temporal order arises only through correlations between observables. The surface elevation of a wave changes with respect to the reading of some other observable (clock), which in our case we choose as related to the vertical velocity gradient. that measures the relevant changes of the dynamics in relation to steepening as observed at a moving point $Q$ on the surface:
@@
where $t_{0}$ is a convenient initial time after which $U_{z}\geq 0$, ensuring that $\tau$ increases monotonically (see bottom-right panel of Fig. 1, note the smoothness of $\tau(t)$ as a result of the integration). This definition applies to waves propagating from left to right along the $x$-axis, where $U_{z}>0$ near the crest of steep waves. [^3] For waves propagating in the opposite direction (right to left), $U_{z}<0$ and one sets $\frac{d\tau}{dt}=-U_{z}$ to maintain a monotonically increasing $\tau$. As waves steepen, the horizontal velocity gradient $U_{z}$ measured at $Q$ grows, eventually becoming very large and formally diverging at the onset of breaking, for instance during the formation of jets associated with spilling. In other words, the internal clock associated with the crest-steepening dynamics ticks more quickly than the external time we use to observe the waves. This motivates defining near breaking, the small parameter
@@
Both parameters $\beta,\epsilon$ are small near breaking, and $\beta\sim O(\epsilon)$, estimated from numerical simulations (Barthelemy et al., 2018). For very steep wave crests near the breaking inception, where $U_{z}\gg 1$ ($\epsilon\ll 1$), our analysis indicates that the focusing parameter $A$ in Eq. (17) is bounded and of order $O(1)$. Near breaking, this implies that $\eta_{xx}$, proportional to surface curvature, grows as fast as $O(\epsilon^{-2})$ to keep $A$ finite. Here, the notation $O(\epsilon^{n})$ denotes terms of order $\epsilon^{n}$ or higher.
@@
We demonstrate that this timescale separation occurs when the special moving point $Q$ is chosen to coincide with the energetic hotspot (Boettger et al., 2023, 2024) (for details, see C). It captures the essential dynamics near the onset of breaking and it reveals a robust fold geometry that guides the evolution of steep waves near breaking.
Our analysis is grounded in first principles and supported by numerical simulations of nonlinear wave groups by Barthelemy et al. (2018). The same separation could not be conclusively established when following other surface points, such as the inflection point (McAllister et al., 2023).
Figure 1: Top panel: snapshots of the (black) surface elevation $\eta(x,t)$ and (red) kinetic energy $\tilde{k}_{e}(x,t)$ of the largest non-breaking wave (labeled 22) in the simulated right-going wave group $C3N5A0.514$ (Barthelemy et al., 2018) at the pre-focusing ($P$), focusing ($F$), and post-focusing ($R$) instants, in the middle top panel of Fig. 3. The energetic hotspot (red square) is located at the front of the wave before focusing and moves to the back after passing through the crest (blue square). The inflection point is also depicted (blue diamond). Bottom panel: (left) time profiles of the horizontal velocity gradient $U_{z}$ tracked at the energetic hotspot of the four major successive waves of the simulated right-going group $C3N5A0.514$ (Barthelemy et al., 2018). These are labeled as $0,1,2,3$ with varying thickness to denote the increasing amplitude, and wave $3$ is breaking; (right) relational fast time $\tau$ as function of the absolute time $t$.
@@ 4 Following the energetic hotspot: the slow-fast model
evaluated at the free surface $z=\eta(x,t)$, where $k_{e}=(u^{2}+w^{2})/2$ is the kinetic energy density of the fluid (see A.4). Drawing on Boettger et al. (2023), we track the most energetic point $Q$ on the wave surface, defined as the location where $\tilde{k}_{e}$ attains its maximum, that is where the spatial kinetic energy gradient $\partial_{x}\tilde{k}_{e}=0$. This energetic point can be regarded as the crest of the kinetic energy surface  $\tilde{k}_{e}(x,t)$. The speed $C_{e}$ of the energetic hotspot reduces to the group velocity for linear wave packets.
To highlight the relating motion of the wave surface crest and energetic point, or hotspot, we consider the simulated breaking wave group $C3N5A0.514$ reported in Barthelemy et al. (2018). The four major successive waves of the group are labeled in order of increasing amplitude, with wave $3$ undergoing breaking. The evolution of the second-largest wave (wave $2$) and its kinetic energy surface is shown in the top panel of Fig. 1. As the wave grows in amplitude (point $P$, pre-focusing phase), the energy crest is located on the front face of the wave, slightly behind the surface crest. At focus (point $F$), where the wave attains its maximum height, the surface crest and the energetic point coincide at the same location. After focus (point $R$), the energetic point shifts to the back face of the wave, behind the surface crest. The other waves in the group have similar evolution, except wave $3$ that breaks near the focusing point (see (Barthelemy et al., 2018) for more details).
During the growth phase, the surface crest slows down (Fedele et al., 2020; Banner et al., 2014; Fedele, 2014a), while the energetic point speeds up. However, the acceleration of the energetic point is insufficient to prevent the surface crest from overtaking it. After focus, the surface crest accelerates, whereas the energetic hotspot decelerates. The corresponding crest and energetic point speeds are shown in the bottom-right panel of Fig. 1.
@@ 4.1 The slow-fast model
The energy gradient $\partial_{x}\tilde{k}_{e}$ vanishes at the energetic hotspot $Q$. Thus, Eq. (69) in A.4) yields
@@
In C, we proved that a slow-fast regime occurs near breaking around the maximum $A=1+O(\epsilon)$, and along the path $\Gamma_{2}$ in the $(m,s)$ plane where $m_{2}\sim(m+s)=0$ (see Eq. (35)). This is where the vertical transport of horizontal momentum density dominates the horizontal advection during focusing when a wave reaches its maximum height, or breaks (see Eq. (16)). Numerical simulations confirm this trend, as illustrated in Figure 4, and is further discussed in section 4.4.
@@
Then, $dC_{e}/d\tau=dC_{r}/d\tau\sim O(\epsilon^{3/2})$. Using these scalings, in C.1 a closed slow-fast system is derived in the variables $(m,A)$ (see Eq. (108)),
@@
Such slow-fast timescale separation is observed in numerical simulations of nonlinear wave groups (Barthelemy et al., 2018; Boettger, 2024). The top panels of Fig. 2 depict the time profiles of the wave parameters $m$, $s$ and $R=m+s$ tracked at the energetic hotspot of the four major successive waves of the simulated right-going group $C3N5A0.514$ (Barthelemy et al., 2018). These are labeled in order of increasing amplitude, with wave $3$ breaking. Near breaking, $R=m+s$ varies adiabatically, more slowly than the surface slope $m$ and the kinematic parameter $s$ over the fast time $\tau$. Similarly, the focusing parameter $A$ and $\beta=C_{e}/U-1$ are also slow variables (see bottom panels of Fig. 2. Thus, the slow variable $A$ acts as an adiabatic forcing for the fast geometric variable $m$.
@@
Figure 2: Top panels: time profiles of the wave parameters $m$, $s$ and $R=m+s$ tracked at the energetic hotspot of the four major successive waves of the simulated right-going group $C3N5A0.514$ (Barthelemy et al., 2018). Bottom panels: same for the slow parameters AA and $\beta=C_{e}/U-1$ tracked at the energetic hotspot. Waves are labeled as $0,1,2,3$ with varying thickness to denote the increasing amplitude, and with wave $3$ breaking. Characteristic slopes are also shown (red lines) and quantified in terms of $\epsilon=1/U_{z}\approx 0.1$.
Figure 3: Fold catastrophe for inviscid and irrotational waves in deep water. (Left) fold boundary $\partial\mathcal{F}_{0}$ in the $(m,A)$ plane and trajectories (solid lines) of the four major successive waves of the simulated right-going wave group $C3N5A0.514$ (Barthelemy et al., 2018). These are labeled as $0,1,2,3$ at the beginning (early time) of the trajectories, and depicted with increasing thickness. The thickest (dark red) line indicates the largest breaking wave. Points $P$, $F$, and $R$ mark the pre-focusing, focusing, and post-focusing phases of the largest non-breaking wave ($2$) (see Fig.1). Theoretical predictions (dashed lines) from the slow-fast model are also depicted. The path of a linear packet is also reported (amplitude $\mu=ka=0.1$ and spectral bandwidth $\nu=0.3$). (Left) the desymmetrized half-domain $m\leq 0$.
@@ 4.2 Fold Catastrophe and Breaking Onset
The fold boundary $\partial\mathcal{F}_{0}$ is shown in the left panel of Fig. 3 and the stable and unstable manifolds of equilibria are depicted with arrows. Equilibria with negative (positive) slope are stable (unstable). These equilibria coalesce at the tip $O$ at $(A=1,m=0)$, forming a marginal state with slope $m=0$. Thus, $\partial\mathcal{F}_{0}$ is normally hyperbolic, except at $O$. This boundary is also symmetric with respect to $m$, reflecting the parity symmetry of the Euler equations in deep water, which allows the same wave to propagate in either direction. In such cases, only the horizontal velocity $U$, its gradient $U_{z}$ and the hotspot speed $C_{e}$ change sign, leading to a corresponding sign change in both $s$ and the slope $m$, except the focusing parameter $A$. The symmetry can be removed by applying symmetry reduction techniques described in Schneider et al. (2004) (see section 4.9 and right panel of Fig. 3 that displays the desymmetrized domain).
The point $O$ is the marginal point of a fold catastrophe (Thom, 1975; Zeeman, 1976; Arnold, 1992). Indeed, we can rewrite the surface slope equation as the gradient flow
@@
defines a cubic fold surface (Thom, 1975; Arnold, 1992).
The right panel of Fig. 3 illustrates the trajectories (blue lines) in the plane $(m,A)$ of the four major successive waves in the simulated right-going wave group $C3N5A0.514$ reported by Barthelemy et al. (2018). These are labeled as $0,1,2,3$ at the beginning (early time) of the trajectories. Line thickness increases with wave amplitude, with the thickest trajectory corresponding to the breaking wave ($3$). Points $P$, $Q$, and $R$ denote the pre-focusing, focusing, and post-focusing phases of the largest non-breaking wave ($2$), whose profiles are shown in Fig. 1. For comparison, the trajectory of a linear Gaussian packet is also included (wave steepness $\mu=ka=0.1$, spectral bandwidth $\nu=0.3$).
All trajectories of the four nonlinear waves, during their growing phase, are attracted toward the point $O$ along the stable manifold of the negative-slope equilibria, and the energetic hotspot lies just behind the wave crest. They follow the parabolic arc toward the point $O$, where the wave crest and the energetic hotspot coincide. The breaking wave ($3$) is the only one that crosses the boundary near $O$, thereby entering its interior and undergoing finite-time blowup, indicating the onset of breaking. The numerical simulations, however, could not resolve the dynamics in the immediate vicinity of, or beyond, breaking. In contrast, the trajectories of the other non-breaking waves are repelled along the unstable manifold of the positive-slope equilibria, leading to the defocusing phase of the crests. A similar behavior is observed in other simulated breaking wave groups (Barthelemy et al., 2018).